During man life of a radioactive element, the fraction that disintegrates is..........


1) e


2) 1/e


3) (e-1)/e


4) e/(e-1)


 


Pls explain........

Grade:12

1 Answers

aman gupta
13 years ago

initial no. of nucleus = N

mean life = 1/λ

after men life = Ne-λt (t=1/λ)

    =Ne-1

    =N/e

fraction left = (N-N/e)/N

ans 3